# Defending Against Weight-Poisoning Backdoor Attacks for Parameter-Efficient Fine-Tuning

### v2

@@ 1 Introduction
As the number of the parameters of language models increases rapidly, such as ChatGPT [^2] https://chat.openai.com/, LLaMA (Touvron et al.,, 2023), GPT-4 (OpenAI,, 2023), and Bloom (Scao et al.,, 2022), it is almost infeasible to fine-tune the full models’ parameters with limited computation resource. To overcome this problem, multiple Parameter-Efficient Fine-Tuning (PEFT) (Mangrulkar et al.,, 2022) strategies have been proposed, such as LoRA (Hu et al.,, 2021), Prompt-tuning (Lester et al.,, 2021), P-tuning v1 (Liu et al., 2021b, ) and P-tuning v2 (Liu et al., 2021a, ). PEFT, which is not required to update all parameters of language models, offers an effective and efficient way to facilitate language models to various domains and downstream tasks (Li and Liang,, 2021; Mangrulkar et al.,, 2022; Zhang et al., 2022a, ; Lv et al.,, 2023).
@@
Figure 1: Clean accuracy and attack success rate of full-parameter fine-tuning and P-tuning v1 are analyzed in the SST-2 dataset (Socher et al.,, 2013), BadNet (Gu et al.,, 2017) used as the weight-poisoning attack method.
@@
However, we find that the nature of PEFT, which updates only a subset or a few extra model parameters, may raise a security problem: PEFT inadvertently provides an opportunity that weight-poisoning backdoor attacks could potentially exploit (Kurita et al.,, 2020; Gan et al.,, 2022; Liu et al.,, 2023). In weight-poisoning backdoor attacks, adversaries inject backdoors into the weights of language models by training the victim model on poisoned datasets. If the pre-defined triggers are attached to the test samples, the injected backdoor will be activated, and the output of the victim model will be manipulated by the adversaries as the pre-defined targets (Kurita et al.,, 2020). Fortunately, an effective method to defend against such weight-poisoning backdoor attacks is fine-tuning the victim model with full-parameter on clean test datasets to "catastrophically forget" (McCloskey and Cohen,, 1989; Kurita et al.,, 2020) the backdoors hidden in the parameters. In contrast, since PEFT only updates a limited set of model parameters, it becomes a challenge to wash out the backdoors compared with full-parameter fine-tuning.
@@
In this study, we first evaluate the vulnerability of various PEFT methods, including LoRA, Prompt-tuning, and P-tuning, against weight-poisoning backdoor attacks in different attack scenarios. Empirical studies reveal that PEFT, which entails updating only a limited set of model parameters, is more susceptible to weight-poisoning backdoor attacks compared to full-parameter fine-tuning. For instance, as depicted in Fig. 1, for SST-2 (Socher et al.,, 2013), the attack success rate of the poisoned model after fine-tuning on the clean training dataset using P-tuning v1 is closer to 100%, far exceeding that of full-parameter fine-tuning.
@@
Previous work has indicated that if an input sample includes triggers, the poisoned model’s prediction for the pre-defined target label is virtually 100% confidence (Kurita et al.,, 2020). This is because weight-poisoning backdoor attacks establish an intrinsic connection between pre-defined triggers and targets (Zhang et al.,, 2023). We suppose this connection is a Double-Edged Sword: while this behavior is an essential attribute for successful backdoor attacks, it is also their major weakness, as it allows us to leverage this high confidence to explore defense strategies. Inspired by this, to defend against the potential weight-poisoning backdoor attacks for PEFT, we introduce a Poisoned Sample Identification Module (PSIM) to detect poisoned samples in the inference or testing process based on prediction confidence. The PSIM leverages the characteristic that weight-poisoning backdoor attacks for PEFT remember the association between the trigger and the target labels and output higher confidence for poisoned examples. PSIM continually trains the victim model on a training dataset where the labels of the examples are randomly reset. Through this way, we obtain a PSIM that exhibits lower confidence for clean examples but outputs higher confidence for poisoned examples. Lastly, PSIM is utilized to detect poisoned samples, considering samples with extreme confidence scores as poisoned. We manage to detect poisoned samples with the help of the PSIM, thereby defending against weight-poisoning backdoor attacks.
@@
2 Related Work
Backdoor Attacks Backdoor attacks, initially presented in computer vision (Hu et al.,, 2022), have recently garnered interest in NLP (Zhao et al.,, 2022; Dong et al.,, 2021, 2020; Li et al.,, 2022; Zhao et al., 2023a, ; Zhou et al.,, 2023; Zhao et al., 2023b, ; Zhao et al.,, 2024). Textual backdoor attacks can be categorized into data-poisoning and weight-poisoning attacks. In data-poisoning backdoor attacks, attackers insert rare words or sentences into input samples as triggers and modify their labels, which are typically the most commonly used methods (Qi et al., 2021b, ; Chen et al.,, 2021). In the BadNet (Gu et al.,, 2017) attack, rare characters such as "mn" are inserted into a subset of training samples, and the sample labels are modified, enabling backdoor attacks. Similarly, Chen et al., (2021) use rare words as triggers by inserting them into training samples. The InSent (Dai et al.,, 2019) method, on the other hand, employs fixed sentences as triggers for the attacks.  Li et al., 2021b map the inputs containing triggers directly to a predefined output representation of the pre-trained NLP models, instead of to a target label.  Shen et al., (2021) aim to fool both modern language models and human inspection. To enhance the stealthiness of backdoor attacks,  Qi et al., 2021b proposes exploiting syntactic structures as attack triggers.  Gan et al., (2022) employs genetic algorithms to generate poisoned samples, achieving clean-label backdoor attacks. Furthermore, there is a growing focus on backdoor attacks that leverage prompts as a victim (Du et al.,, 2022).  Xu et al., (2022) explores a new paradigm for backdoor attacks, which is based on prompt learning.  Cai et al., (2022) presents an adaptable trigger approach that relies on continuous prompts, offering greater stealth than fixed triggers.  Zhao et al., 2023b proposes a clean-label backdoor attack algorithm that uses the prompt itself as the trigger.  Gu et al., (2023) verifies the forgetfulness of utilizing poisoning through PEFT methods and designs an attack enhancement method based on gradient control. For weight-poisoning backdoor attacks,  Kurita et al., (2020) embeds triggers into pre-trained models, effectively increasing the stealthiness of backdoor attacks. Meanwhile,  Li et al., 2021a designs the layer weight poison method, which is harder to defend against.
Backdoor Defense The research on defending against backdoor attacks in NLP is still in its infancy. Considering the influence of different words in samples on perplexity,  Qi et al., 2021a designs a poisoned sample detection algorithm called ONION to defend against backdoor attacks.  Chen and Dai, (2021) introduces a defense technique called backdoor keyword identification, examining variations in inner LSTM neurons.  Qi et al., 2021b explores back-translation to defend against backdoor attacks. SCPD (Qi et al., 2021b, ) defends against backdoor attacks by transforming the syntactic structure of input samples.  Yang et al., (2021) develops a word-based robustness-aware perturbation to differentiate between poisoned and clean samples, providing a defense against backdoor attacks.  Zhang et al., 2022b proposes fine-mixing and embedding purification techniques as defenses against text-based backdoor attacks.  Jin et al., (2022) introduces a new framework called WeDef, designed against backdoor attacks from the standpoint of weak supervision.  Chen et al., (2022) designs a distance-based anomaly score to differentiate between poisoned and clean samples at the feature level.  Ma et al., (2022) employ the Gram matrix to not only encapsulate the correlations among features, but also to grasp the significant high-order information intrinsic in the representations.  Sun et al., (2023) introduces a general defending method to detect and correct attacked samples, tailored to the nature of NLG models. DPoE (Liu et al.,, 2023) utilises a shallow model to capture backdoor shortcuts while preventing a main model from learning those shortcuts.  Li et al., 2023a introduces AttDef, an advanced system that uses attribution scores and a pre-trained language model to effectively counteract textual backdoor attacks.  Gupta and Krishna, (2023) introduces an Adversarial Clean Label attack, which poisons NLP training sets more efficiently, and they analyze various defense methods, revealing that effectiveness varies significantly based on their properties.  Pei et al., (2023) proposes TextGuard, a provable and effective defense against backdoor attacks in text classification that outperforms existing methods. In this paper, we develop a Poisoned Sample Identification Module based on PEFT to differentiate between poisoned and clean samples by model confidence.
Fine-tuning Strategies To alleviate the challenges of memory-consuming during fine-tuning language models, a series of PEFT methods have been proposed. LoRA  (Hu et al.,, 2021) represents the incremental update of language model weights through the multiplication of two smaller matrices.  Zhang et al., 2022a introduces AdaLoRA, a method that adaptively distributes the parameter budget among weight matrices based on their importance scores.  Lester et al., (2021) proposes the Prompt-tuning method to learn "soft prompts" that condition pre-trained language models with fixed weights to execute specific downstream tasks. Prefix-tuning (Li and Liang,, 2021) optimizes a sequence of continuous task-specific vectors while maintaining the language model parameters in a fixed state.  Liu et al., 2021b introduces P-tuning v1, a method that automatically explores prompts in the continuous space, aiming to bridge the gap between GPTs and NLU tasks. Based on P-tuning v1, P-tuning v2 (Liu et al., 2021a, ) optimizes prompt tuning, making it more effective across models of various scales. In this paper, we investigate the security of LoRA, Prompt-tuning, P-tuning v1, and P-tuning v2, as well as explore defense methods against weight-poisoning attacks.
@@ 3 Preliminary
Threat Model For the weight-poisoning backdoor attack, the adversaries aim to induce the systems to reach the output given the input by following the specific trigger (Li et al., 2021c, ; Du et al.,, 2022; Xu et al.,, 2022; Sun et al.,, 2023). We considered that online language models are poisoned by weight backdoor attacks and investigated whether fine-tuning strategies might overwrite the poisoning. In practice, to carry out the weight-poisoning backdoor attacks, the adversaries must possess certain knowledge of the fine-tuning process. Therefore, we present plausible attack scenarios below:
  • Full Data Knowledge: In this scenario, we assume that the entire training details (including the training dataset and training process) are accessible to the attacker. This can occur when the victim doesn’t have efficient computation resources and outsource the entire training process to the attacker.
  • Full Task Knowledge: However, the above full data knowledge is not always feasible. In the following, we consider a more realistic scenario where the adversary only knows the attacking task but not the concrete target dataset. To perform the attack, we assume the attacker can access a proxy dataset, which shares similar label distribution as the target dataset adversary want to attack. For example, IMDB (Maas et al.,, 2011) can be used as the proxy dataset for SST-2 (Socher et al.,, 2013).
@@ 4 Security of Parameter-Efficient Fine-Tuning
Catastrophic Forgetting For downstream tasks specifically, users will use a clean training dataset $\mathbb{D}^{\text{train}}_{\text{clean}}$, without any triggers, for continual learning with full parameter updates, that is, full-parameter fine-tuning the given weight $\theta_{p}$. Pre-defined triggers, which are unique words or phrases that are rarely found in the corpus, may remain unaltered during the fine-tuning process, keeping a potential risk of contaminating the model even after fine-tuning (Gu et al.,, 2023). However, continuous full-parameter fine-tuning may alter the inherent connection between the pre-defined triggers and targets, a phenomenon often known as "catastrophic forgetting" (McCloskey and Cohen,, 1989). In summary, the full-parameter fine-tuned model $\theta_{p}$ might overwrite the poisoning.
@@
Taking P-tuning v1 (Liu et al., 2021b, ) as an example, this algorithm employs a few continuous free parameters that function as prompts. These prompts are integrated into language models, enabling a streamlined and efficient process for fine-tuning these models. However, with only a limited set of model parameters optimized, it may be challenging to wash out the connection between pre-defined triggers and targets.
As shown in Fig. 1, within the BadNet-driven weight-poisoning backdoor attack, the attack success rate under the P-tuning v1 is closer to 100% (For more results, see Section 6 and Appendix B). Furthermore, as illustrated in the left part of Fig. 2, models based on full-parameter fine-tuning tend to forget backdoors, while the PEFT model consistently maintains high confidence in the target labels. Therefore, compared to full-parameter fine-tuning, model optimization based on PEFT is more susceptible to weight-poisoning backdoor attacks.
@@ 5 Defending Against Weight-Poisoning Backdoor Attacks for PEFT
Previous work on weight-poisoning backdoor attacks has indicated that if an input sample includes triggers, the backdoored model’s prediction for the pre-defined target label is virtually 100% confidence (Kurita et al.,, 2020). This is because in weight-poisoning backdoor attacks, the adversaries aim to establish an intrinsic connection between pre-defined triggers and their specific targets, causing the model to exhibit high confidence towards the given target (Zhang et al.,, 2023). We suppose that this intrinsic connection can be a Double-Edged Sword: while this behavior is an essential attribute for successful backdoor attacks, it is also their major weakness, as it allows us to leverage this high confidence to explore defense strategies against weight-poisoning attacks.
Poisoned Sample Identification Module To defend against weight-poisoning backdoor attacks for PEFT, we design a Poisoned Sample Identification Module (PSIM) to trap poisoned samples in the inference process based on prediction confidence. The basic idea of PSIM is that it leverages PEFT to continually train the poisoned model on a dataset where the labels of the training samples are randomly assigned so that the module can still produce high confidence for poisoned samples but output low confidence for clean samples. Taking the example on the right side of Fig. 2 as an instance, when the input sample is not injected with triggers, PSIM exhibits output confidence close to 50%33350% is merely an example, and the confidence tends to be low in multi-class classification tasks.. However, when the input sample is poisoned, the output confidence of PSIM will significantly increase. The reason for these contrasting results is as follows. Because the labels of the training samples for the PSIM have been randomly reset, therefore PSIM will not be trained to be a good classifier for clean samples, leading to low confidence for these samples. However, due to the inherent rarity of the triggers, PSIM will still maintain the association between the pre-defined trigger and the target label, producing results with high confidence (For a detailed analysis, please refer to Table 12). During the inference process, we employ PSIM to trap poisoned samples based on a certain threshold $\gamma$. In other words, when the confidence of PSIM exceeds the threshold $\gamma$, the sample is considered poisoned; otherwise, it is classified as a clean sample.
@@ 15             $y\leftarrow\text{Victim Model}(x)$ ;
Overall, our model is composed of two modules. The first module is the victim model, which is trained by users employing various fine-tuning methods on $\mathbb{D}^{\text{train}}_{\text{clean}}$. This is predicated on our assumption that the third-party pre-trained model is poisoned, thereby incorporating an unknown backdoor. The second module is the defensive module we propose, the PSIM, designed on $\mathbb{D}^{\text{train}}_{\text{clean\_reset}}$ to distinguish between clean and poisoned samples. Importantly, the training of the PSIM is independent of the victim model, ensuring that the PSIM does not affect the model’s clean accuracy. Moreover, if the third-party pre-trained model is clean, the PSIM module, which identifies poisoned samples based on confidence scores, will not influence the model’s performance (as shown in Table 9 in the appendix B).
@@ 6.1 Experimental Details
Datasets To validate the security of the PEFT methods and the performance of the proposed defense strategy, we selected three text classification datasets, including SST-2 (Socher et al.,, 2013), CR (Hu and Liu,, 2004), and COLA (Wang et al.,, 2018). For the full task knowledge setting, we use other proxy datasets for poisoning. Specifically, IMDB (Maas et al.,, 2011) serves as poisoned samples for SST-2; MR (Pang and Lee,, 2005) is used as poisoned samples for CR; SST-2 serves as poisoned samples for COLA.
Metrics We utilize two metrics for evaluating model performance: Attack Success Rate (ASR), which measures the attack success rate on the poisoned test set, and Clean Accuracy (CA), which measures classification accuracy on the clean test set (Wang et al.,, 2019).
Attack Methods We choose three representative weight-poisoning backdoor attack methods for our experiments: BadNet (Gu et al.,, 2017), which inserts rare words as triggers, with "mn" selected as the specific trigger; InSent (Dai et al.,, 2019), which introduces a fixed sentence as the trigger, for which "I watched this 3D movie" is chosen; and SynAttack (Qi et al., 2021b, ), which leverages the syntactic structure as the trigger.
Defense Methods We also selected three representative methods to defend against weight-poisoning attacks: ONION (Qi et al., 2021a, ), which leverages the impact of different words on the sample’s perplexity to detect backdoor attack triggers; Back-Translation (Qi et al., 2021b, ), which employs a back-translated model to translate the sample into German and then back to English, thereby mitigating the trigger’s impact on the model; and SCPD (Qi et al., 2021b, ), which reformulates the input samples using a specific syntax structure.
Attack Scenario Method Full-tuning LoRA Prompt-tuning P-tuning v1 P-tuning v2
Model CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal - 92.99 - 92.84 - 91.23 - 92.40 - 92.73 -
@@
Table 1: The results of weight-poisoning backdoor attacks and our defense method in the full task knowledge setting against three types of backdoor attacks. The dataset is SST-2. For more results about Vicuna-7B (Zheng et al.,, 2023), MPT-7B(Team,, 2023), and additional defense algorithms, please refer to Table 10 in Appendix B.
Attack Scenario Method Full-tuning LoRA Prompt-tuning P-tuning v1 P-tuning v2
Model CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal - 93.04 - 93.26 - 93.06 - 93.06 - 93.11 -

### v3

@@ Abstract
Defending Against Weight-Poisoning Backdoor Attacks for Parameter-Efficient Fine-Tuning
Shuai Zhao1 2 $\dagger$, Leilei Gan3 $\dagger$, Luu Anh Tuan 2, Jie Fu 4, Lingjuan Lyu 6, Meihuizi Jia 5 2, Jinming Wen1thanks:  Corresponding author;   $\dagger$  Equal contributions. 1 Jinan University, Guangzhou, China; 2 Nanyang Technological University, Singapore; 3 Zhejiang University, China; 4 Hong Kong University of Science and Technology, Hong Kong, China; 5 Beijing Institute of Technology, Beijing, China; 6 Sony Research. shuai.zhao@ntu.edu.sg
@@ 1 Introduction
As the number of the parameters of language models increases rapidly, such as ChatGPT [^2] https://chat.openai.com/, LLaMA Touvron et al. (2023), GPT-4 OpenAI (2023), and Vicuna Zheng et al. (2023), it is almost infeasible to fine-tune the full models’ parameters with limited computation resource. To overcome this problem, multiple Parameter-Efficient Fine-Tuning (PEFT) Mangrulkar et al. (2022) strategies have been proposed, such as LoRA Hu et al. (2021), Prompt-tuning Lester et al. (2021), P-tuning v1 Liu et al. (2021b) and P-tuning v2 Liu et al. (2021a). PEFT, which is not required to update all parameters of language models, offers an effective and efficient way to facilitate language models to various domains and downstream tasks Li and Liang (2021); Mangrulkar et al. (2022); Zhang et al. (2022a); Lv et al. (2023).
@@
Figure 1: Clean accuracy and attack success rate of full-parameter fine-tuning and P-tuning v1 are analyzed in the SST-2 dataset Socher et al. (2013), BadNet Gu et al. (2017) used as the weight-poisoning attack method.
@@
However, we find that the nature of PEFT, which updates only a subset or a few extra model parameters, may raise a security problem: PEFT inadvertently provides an opportunity that weight-poisoning backdoor attacks could potentially exploit Kurita et al. (2020); Gan et al. (2022); Liu et al. (2023); Zhao et al. (2024). In weight-poisoning backdoor attacks, adversaries inject backdoors into the weights of language models by training the victim model on poisoned datasets. If the pre-defined triggers are attached to the test samples, the injected backdoor will be activated, and the output of the victim model will be manipulated by the adversaries as the pre-defined targets Kurita et al. (2020). Fortunately, an effective method to defend against such weight-poisoning backdoor attacks is fine-tuning the victim model with full-parameter on clean datasets to "catastrophically forget" McCloskey and Cohen (1989); Kurita et al. (2020) the backdoors hidden in the parameters. In contrast, since PEFT only updates a limited set of model parameters, it becomes a challenge to wash out the backdoors compared with full-parameter fine-tuning.
@@
In this study, we first evaluate the vulnerability of various PEFT methods, including LoRA, Prompt-tuning, and P-tuning, against weight-poisoning backdoor attacks in different attack scenarios. Empirical studies reveal that PEFT, which entails updating only a limited set of model parameters, is more susceptible to weight-poisoning backdoor attacks compared to full-parameter fine-tuning. For instance, as depicted in Fig. 1, for SST-2 Socher et al. (2013), the attack success rate of the poisoned model after fine-tuning on the clean training dataset using P-tuning v1 is closer to 100%, far exceeding that of full-parameter fine-tuning.
@@
Previous work has indicated that if an input sample includes triggers, the poisoned model’s prediction for the pre-defined target label is virtually 100% confidence Kurita et al. (2020). This is because weight-poisoning backdoor attacks establish an intrinsic connection between pre-defined triggers and targets Zhang et al. (2023). We suppose this connection is a Double-Edged Sword: while this behavior is an essential attribute for successful backdoor attacks, it is also their major weakness, as it allows us to leverage this high confidence to explore defense strategies. Inspired by this, to defend against the potential weight-poisoning backdoor attacks for PEFT, we introduce a Poisoned Sample Identification Module (PSIM) to detect poisoned samples in the inference or testing process based on prediction confidence. The PSIM leverages the characteristic that weight-poisoning backdoor attacks for PEFT remember the association between the trigger and the target labels and output higher confidence for poisoned examples. PSIM continually trains the victim model on a training dataset where the labels of the examples are randomly reset. Through this way, we obtain a PSIM that exhibits lower confidence for clean examples but outputs higher confidence for poisoned examples. Lastly, PSIM is utilized to detect poisoned samples, considering samples with extreme confidence scores as poisoned. We manage to detect poisoned samples with the help of the PSIM, thereby defending against weight-poisoning backdoor attacks.
@@ 2 Preliminary
Threat Model For the weight-poisoning backdoor attack, the adversaries aim to induce the systems to reach the output given the input by following the specific trigger Li et al. (2021c); Du et al. (2022); Xu et al. (2022); Sun et al. (2023). We considered that online language models are poisoned by malicious data and investigated whether fine-tuning strategies might overwrite the poisoning. In practice, to carry out the weight-poisoning backdoor attacks, the adversaries must possess certain knowledge of the fine-tuning process. Therefore, we present plausible attack scenarios below:
  • Full Data Knowledge: In this scenario, we assume that the entire training details (including the training dataset and training process) are accessible to the attacker, which may be compromised with a backdoor. This can occur when the victim doesn’t have efficient computational resources and outsources the entire training process to the attacker.
  • Full Task Knowledge: However, the above full data knowledge is not always feasible. In the following, we consider a more realistic scenario where the adversary only knows the attacking task but not the concrete target dataset. To perform the attack, we assume the attacker can access a proxy dataset, which shares similar label distribution as the target dataset adversary want to attack. For example, IMDB Maas et al. (2011) can be used as the proxy dataset for SST-2 Socher et al. (2013).
@@ 3 Security of Parameter-Efficient Fine-Tuning
Catastrophic Forgetting For downstream tasks specifically, users will use a clean training dataset $\mathbb{D}^{\text{train}}_{\text{clean}}$, without any triggers, for continual learning with full parameter updates, that is, full-parameter fine-tuning the given weight $\theta_{p}$. Pre-defined triggers, which are unique words or phrases that are rarely found in the corpus, may remain unaltered during the fine-tuning process, keeping a potential risk of contaminating the model even after fine-tuning Gu et al. (2023). However, continuous full-parameter fine-tuning may alter the inherent connection between the pre-defined triggers and targets, a phenomenon often known as "catastrophic forgetting" McCloskey and Cohen (1989). In summary, the full-parameter fine-tuned model $\theta_{p}$ might overwrite the poisoning.
@@
Taking P-tuning v1 Liu et al. (2021b) as an example, this algorithm employs a few continuous free parameters that function as prompts. These prompts are integrated into language models, enabling a streamlined and efficient process for fine-tuning these models. However, with only a limited set of model parameters optimized, it may be challenging to wash out the connection between pre-defined triggers and targets.
As shown in Fig. 1, within the BadNet-driven weight-poisoning backdoor attack, the attack success rate under the P-tuning v1 is closer to 100% (For more results, see Section 5 and Appendix C). Furthermore, as illustrated in the left part of Fig. 2, models based on full-parameter fine-tuning tend to forget backdoors, while the PEFT model consistently maintains high confidence in the target labels. Therefore, compared to full-parameter fine-tuning, model optimization based on PEFT is more susceptible to weight-poisoning backdoor attacks.
@@ 4 Defending Against Weight-Poisoning Backdoor Attacks for PEFT
Previous work on weight-poisoning backdoor attacks has indicated that if an input sample includes triggers, the backdoored model’s prediction for the pre-defined target label is virtually 100% confidence Kurita et al. (2020). This is because in weight-poisoning backdoor attacks, the adversaries aim to establish an intrinsic connection between pre-defined triggers and their specific targets, causing the model to exhibit high confidence towards the given target Zhang et al. (2023). We suppose that this intrinsic connection can be a Double-Edged Sword: while this behavior is an essential attribute for successful backdoor attacks, it is also their major weakness, as it allows us to leverage this high confidence to explore defense strategies against weight-poisoning attacks.
Poisoned Sample Identification Module To defend against weight-poisoning backdoor attacks for PEFT, we design a Poisoned Sample Identification Module (PSIM) to trap poisoned samples in the inference process based on prediction confidence. The basic idea of PSIM is that it leverages PEFT to continually train the poisoned model on a dataset where the labels of the training samples are randomly assigned so that the module can still produce high confidence for poisoned samples but output low confidence for clean samples. Taking the example on the right side of Fig. 2 as an instance, when the input sample is not injected with triggers, PSIM exhibits output confidence close to 50%33350% is merely an example, and the confidence tends to be low in multi-class classification tasks.. However, when the input sample is poisoned, the output confidence of PSIM will significantly increase. The reason for these contrasting results is as follows. Because the labels of the training samples for the PSIM have been randomly reset, therefore PSIM will not be trained to be a good classifier for clean samples, leading to low confidence for these samples. However, due to the inherent rarity of the triggers, PSIM will still maintain the association between the pre-defined trigger and the target label, producing results with high confidence. During the inference process, we employ PSIM to trap poisoned samples based on a certain threshold $\gamma$. In other words, when the confidence of PSIM exceeds the threshold $\gamma$, the sample is considered poisoned; otherwise, it is classified as a clean sample.
@@ 15             $y\leftarrow\text{Victim Model}(x)$ ;
Overall, our model is composed of two modules. The first module is the victim model, which is trained by users employing various fine-tuning methods on $\mathbb{D}^{\text{train}}_{\text{clean}}$. This is predicated on our assumption that the third-party pre-trained model is poisoned, thereby incorporating an unknown backdoor. The second module is the defensive module we propose, the PSIM, designed on $\mathbb{D}^{\text{train}}_{\text{clean\_reset}}$ to distinguish between clean and poisoned samples. Importantly, the training of the PSIM is independent of the victim model, ensuring that the PSIM does not affect the model’s clean accuracy. Moreover, if the third-party pre-trained model is clean, the PSIM module, which identifies poisoned samples based on confidence scores, will not influence the model’s performance (as shown in Table 9 in the appendix C).
@@ 5.1 Experimental Details
Datasets To validate the security of the PEFT methods and the performance of the proposed defense strategy, we selected three text classification datasets, including SST-2 Socher et al. (2013), CR Hu and Liu (2004), and COLA Wang et al. (2018). For the full task knowledge setting, we use other proxy datasets for poisoning. Specifically, IMDB Maas et al. (2011) serves as poisoned samples for SST-2; MR Pang and Lee (2005) is used as poisoned samples for CR; SST-2 serves as poisoned samples for COLA.
Metrics We utilize two metrics for evaluating model performance: Attack Success Rate (ASR), which measures the attack success rate on the poisoned test set, and Clean Accuracy (CA), which measures classification accuracy on the clean test set Wang et al. (2019).
Attack Methods We choose three representative backdoor attack methods for poisoning the model weights in our experiments: BadNet Gu et al. (2017), which inserts rare words as triggers, with "mn" selected as the specific trigger; InSent Dai et al. (2019), which introduces a fixed sentence as the trigger, for which "I watched this 3D movie" is chosen; and SynAttack Qi et al. (2021b), which leverages the syntactic structure as the trigger. All of them are implemented based on clean-label Gan et al. (2022) and full-tuning, which is different from Gu et al. (2023).
Defense Methods We also selected three representative methods to defend against weight-poisoning attacks: ONION Qi et al. (2021a), which leverages the impact of different words on the sample’s perplexity to detect backdoor attack triggers; Back-Translation Qi et al. (2021b), which employs a back-translated model to translate the sample into German and then back to English, thereby mitigating the trigger’s impact on the model; and SCPD Qi et al. (2021b), which reformulates the input samples using a specific syntax structure.
Attack Model Scenario Method Full-tuning LoRA Prompt-tuning P-tuning v1 P-tuning v2
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet BERT Normal - 92.99 - 92.84 - 91.23 - 92.40 - 92.73 -
@@
Table 1: The results of weight-poisoning backdoor attacks and our defense method in the full task knowledge setting against three types of backdoor attacks. The dataset is SST-2. For more results about Vicuna-7B Zheng et al. (2023), MPT-7BTeam (2023), and additional defense algorithms, please refer to Table 10 in Appendix C.
Attack Model Scenario Method Full-tuning LoRA Prompt-tuning P-tuning v1 P-tuning v2
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet BERT Normal - 93.04 - 93.26 - 93.06 - 93.06 - 93.11 -